OpenStudy (anonymous):

an argon ion emits light at 532 nm. What is the frequency of this radiation? is it visible to the human eye? if so, predict the color associated with this wavelength

OpenStudy (chmvijay):

frequency * wavelength = C frequency =c/wavelength where C is velocity of light in vacuum find frequency wavelength is already given 532 nm if the wavelength falls in visible region(400-700 nm) then u can see it VIBGYOR find color from this Violet is 400nm and red is 700nm

OpenStudy (anonymous):

That's my attachment to you. I hope it helps. Thank you!

OpenStudy (whpalmer4):

@rea1304 you didn't do that quite correctly. \[\lambda = 532\text{ nm} * 1 \text{ m}/10^9\text{ nm} = 5.32*10^{-7}\text{ m}\] \[c = f \lambda\]\[f = \frac{c}{\lambda} = \frac{2.99792*10^8\text{ m/s}}{5.32*10^{-7}\text{ m}} = 5.63519*10^{14} \text{ s}^{-1} = 563.519 \text{ THz} \] Always look at the answer and see if it makes sense. Your answer was \(5.64* 10^{-5} \text{ Hz}\) — that's a vibration every 4.9 hours :-)